OpenStudy (anonymous):

Factor completely −7x2 + 14x − 21.

OpenStudy (anonymous):

no

OpenStudy (abdullahm):

First factor out -7 from it

OpenStudy (anonymous):

\[-7x^2+14-21=0\]\[-x^2+2x-3=0\]\[x^2-2x+3=0\]

OpenStudy (anonymous):

now continue from there

OpenStudy (anonymous):

thank u @saseal

OpenStudy (abdullahm):

you're supposed to guide her...

OpenStudy (anonymous):

you need explanation?

OpenStudy (abdullahm):

not give her steps on how to do the tough parts and give her the easy part to solve...

OpenStudy (anonymous):

-7 * (x^2 - 2x +3 ) got it

OpenStudy (anonymous):

no its not even close yet

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now complete the square for the chunk inside the parentheses

OpenStudy (anonymous):

see what ya get